|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
|
|
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
|
|
SASE1 Status and SASE3 Readiness
Technical integration and coordination
|
| Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 292
Entry time: 16 Jun 2017, 10:18
|
| Author: |
Monica Aresté |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Status and plans after SASE1 Readiness Meeting |
|
|
General:
- E. Boyd update the status about the tasks for radiation measurements the 26th June. See presentation attached.
- J. Gruenert suggest to use similar interface as is being use for the machine and will send the link to G. Wellenreuther.
XTDs
- Images for Photo commissioning on the tunnel are stored by a temporary solution. J. Grunert request when the DAS routines will work, CAS will prepare a short presentation for the meeting next Friday to show the progress on that issue.
- This weekend is scheduled to commissioning with a WL of 1,5Å and 30 bunches into the dump for first time.
FXE
- Radiation safety measures done last Wednesday.
- Leaks in the Vacuum system already fixed.
- Cables for motors already repaired by JOWO this week. Reparations will continue next week after urgent cables for SPB are done.
- In general loops are working but there are still some minor issues, AE working in some loops updates.
- Meting on Monday between FXE and AE to define Interlock to deploy.
SPB
- Loop 1 is working, AE to fix some mismatchs in loops 2 and 3.
- Export for loop 7 is done, AE is working to generate the firmware. EET has two persons working in the exports for loops 4 and 12.
- Ion pump cables repaired and tests on going.
- All connector reparations in the OPT-hutch are done and the test of the motors are on going.
- JOWO will keep working in 20 urgent cables for SPB and then move back to FXE.
Detector
- AGIPD is fully powered, there are still some issues for the Intelock that needs AE support.
- More details see attached file provided by Detector Group.
Laser
- The humidity in the hutch still oscillates, this shoud be fixed urgently since it is preventing further commissioning.
- Image processor not working since each 2-3 hours stops, CAS to fix that issue.
- There are motors not working, CAS to investigate.
Electronics
- First priority is is to have the Vacuum system for FXE and SPB operational at the end of June. Next two weeks AE will be mainly focus in the loops that are still missing (SPB 4, 5, 7,12) or still not running (SPB 2, 3)
- Elmo controllers not integrated, AE will test a solution in SPB on next week. If it works will be deployed also for the tunnels but this is not first priority and will be done in July.
Software
- CAS is working to have a more robust system, S. Brockhauser point out that new request from instruments should we send to CAS to define priorities.
- To prepare a short presentation for next Friday about the issue of the tunnel DAS to store images (see point in XTDs)
|
|
|